OpenStudy (anonymous):

find the two numbers whose differences is 4 and the difference of their squares is 88

Directrix (directrix):

@luhan Let the numbers be x and y. x - y = 4 because the two numbers differ by 4 -------- the difference of their squares is 88 x^2 - y^2 = 88 ----------------- Then, solve these two equations simultaneously: x - y = 4 x^2 - y^2 = 88 ------------------ If you are allowed to use technology in your work, go here: http://www.wolframalpha.com/input/?i=%28x+-+y+%3D+4%29+and+%28x%5E2+-+y%5E2+%3D+88%29 If not, use substitution to crank out the answers by hand.